User Tools

Site Tools


lbaops:lbaoct2016:gg081bcdlog

The C-band VLBI scans between 0840 and 0900 have been changed to K-band, as we cannot support rapid receiver changes.

Recording using the DBBC/Mark5C.

Fringes found, although very weak. Problem appears to be a serious pointing offset (~1.8' in Azimuth), unable to be easily corrected.

Main experiment uses a higher IF power target than in fringe testing.

Pols correct.

Recording to HOB+1009/16000/1024

Good weather, system temperature appears very good.

lbaops/lbaoct2016/gg081bcdlog.txt · Last modified: 2016/10/18 18:40 by breid